# GCLS2: Towards Efficient Community Detection using Graph Contrastive Learning with Structure Semantics

### v1

@@ IV-D Structure Contrastive Learning
1Input: i) the data graph $G$ with adjacency matrix $A$ and attribute matrix $X$, and ii) the number of training epochs $\varepsilon$
@@
1:  $G^{H}\leftarrow G$, $A^{H}\leftarrow A$
22:  $S\leftarrow$ Eq. 1 with ($G^{H}$)
3:  for epoch $\leftarrow 1,2,\cdots,\varepsilon$ do
4:     $S^{\prime}\leftarrow$ Eq. 2 with ($S$)
5:     $X^{\prime}\leftarrow$ Eq. 3 with ($X$)
6:     $Z\leftarrow$ Eq. 4 with ($A,AGG(S^{\prime},X^{\prime})$)
7:     $Z^{H}\leftarrow$ Eq. 7 with ($A^{H},AGG(S^{\prime},X^{\prime})$)
8:     $\mathcal{L}\leftarrow$ Eq. 11 with ($Z,Z^{H}$)
9:     repeat steps 4-8 to update the network parameters by gradient descent to minimize $\mathcal{L}$
10:  end for
11:  return  $Z$
@@
The detailed GCLS2 training process is shown in Algorithm 1. Specifically, we first obtain the high-level structure graph $G^{H}$ and high-level adjacency matrix $A^{H}$ from the original data graph $G$, and then extract the structure similarity matrix $S$ from $G^{H}$ (lines 1-2). For each training iteration, we first use the SSS module to extract structural semantics $S^{\prime}$ and attribute semantics $X^{\prime}$ (lines 3-5). Next, we employ a GCN with weight shared to extract node features from both $G$ and $G^{H}$, resulting in the node feature matrix $Z$ and the high-level node feature matrix $Z^{H}$ (lines 6-7). After that, we adjust the network parameters by minimizing structure contrastive learning $\mathcal{L}$ on $Z$ and $Z^{H}$ (lines 8-9). Finally, after $\varepsilon$ training iterations, we return the node features $Z$ obtained for subsequent community detection (lines 10-11).
TABLE I: Statistics of the tested real-world graph datasets. The “#Attrs.” denotes the number of attributes, the “NA” denotes no attributes, and the “#Comms.” denotes the number of communities.
Datasets #Nodes #Edges Avg. Deg. #Attrs. #Comms.
Cora 2,708 5,429 4.01 1,433 7
Citeseer 3,327 4,732 2.84 3,703 6
PubMed 19,717 44,338 4.50 500 3
CoauthorCS 18,333 81,894 8.93 6,805 15
AmazonPhotos 7,650 119,081 31.13 745 8
Email-Eu 1,005 25,571 50.93 NA 42
@@
V Experiments
In this section, we reported the effectiveness of the test results of our proposed GCLS2 on real-world datasets by answering the following three research questions.
@@ V Experiments
RQ2 (Applicability of GCLS2): What are the impacts of different parameter settings (e.g., semantic dimension $d$ and temperature parameter $\tau$) on the performance of GCLS2?
RQ3 (Benefits of GCLS2): Can GCLS2 significantly improve the accuracy and modularity of community detection?
V-A Experimental Settings
Datasets and Baselines: We use six public real-world graph datasets: the widely-used citation networks (i.e., Cora, Citeseer, and PubMed [29]), the co-authorship network and the product co-purchase network (i.e., CoauthorCS and AmazonPhotos [30], resp.), and the e-mail network between members of a European research institution (Email_Eu [31]) to evaluate the effectiveness of our GCLS2 method on the community detection task. The details of the datasets are in Table I. We compare our approach with one supervised learning baseline (i.e., GCN [18]) and seven unsupervised learning baselines, including DGI [20], MVGRL [21], GRACE [11], GCA [12], SUGRL [32], NCLA [22], and ASP [33].
Metrics: We use classical community detection metrics, normalized mutual information (NMI), accuracy (ACC), and Macro-F1 score (MF1) [34], to evaluate the model detection effectiveness. For all metrics above, the higher their values the better.
@@ V-A Experimental Settings
We ran all the experiments on a PC with an NVIDIA RTX 3090 GPU and 24 GB memory. All algorithms were implemented in Python and executed with Python 3.11 interpreter. The datasets and models are in the supplementary material.
V-B The Performance Evaluation – RQ1
To answer RQ1, we evaluated the performance of the baselines and GCLS2 with and without attributes. For the without-attribute case, we tested by replacing the attribute matrix with an adjacency matrix. Table II shows the performance of baselines and GCLS2 on different graph networks with attributes. Table III shows the performance without attributes. Overall, the tables show that our GCLS2 performs strongly on all six datasets, especially in the case of without attributes, where it has quite an advantage. These performances validated the excellence of our GCLS2 method. Specifically, we had the following observations.
Firstly, GCLS2 has an advantage in comparison to GCN, which we believe is due to the fact that GCLS2 deeply mines the structure information of the community and enhances the representation of this information through structure contrastive learning, and also compared to GCN, which has an average reduction of 4.76% in the case of without attributes, GCLS2 has a reduction of 3.05%, where excluding the Citeseer dataset, it has only a reduction of 1%. This emphasizes the importance of considering structure semantics in community detection.
Secondly, we find that the node-graph level contrastive learning methods (DGI, MVGRL) are significantly more effective than the node-node level (GRACE, NCLA, GCLS2) in the CoauthorCS dataset. We believe that the CoauthorCS with a large number of attributes (i.e., 6,805) is represented on the data graph with rich information, which is more sensitive to the community information, and at the same time, removing the attribute information, the node-graph level effect is not outstanding, which confirms our thoughts. In contrast, our proposed GCLS2, which also focuses on structure semantic representation at the node-node level, has the highest detection performance without attribute information.
V-C The Ablation Study – RQ2
To answer RQ2, we studied the variants of the proposed GCLS2 method and analyzed the performance of each of the important modules of Figure 2. Specifically, we had three variants: i) we separated the structure similarity matrix $S$, which represents dense community information, as a variant, i.e., GCLS2 w/o $S$; ii) we separated the structure similarity semantic (SSS) module, which extracts low-level semantic representations as a variant, i.e., GCLS2 w/o SSS; iii) we separated the structure contrastive loss that reinforces the expression of community structure information as a variant, i.e., GCLS2 w/o SCL, and the rest of the processing for each variant was the same as for GCLS2. Table IV shows the performance of the GCLS2’s variants on Cora and Email-Eu datasets. From Table IV, the GCLS2 consistently has the highest accuracy among all the variants on the two datasets. Compared to the three variants, the GCLS2 achieves significant gains on Email-Eu and Cora. Especially in the Email-Eu, the accuracy of using SCL increased by 8%. This reflects that the structure semantic information is effective in community detection, and the structure contrastive learning has shown good performance improvement on Email-Eu.
TABLE IV: Performance of the variants on Cora and Email-Eu.
@@ V-C The Ablation Study – RQ2
V-D The Hyperparameter Analysis – RQ3
Figure 4 shows the sensitivity analysis of the hyperparameters (structure semantic dimension $d$ and temperature parameter $\tau$) of GCLS2 on Cora and Email-Eu datasets. Then we can answer the RQ3. From the results, the ACC, NMI, and MF1 metrics are stable on Cora and Email-Eu datasets for $d$={8, 16, 32, 64, 128}. However, the temperature parameter $\tau$ has the best performance when the $\tau$=1. Since the $\tau$ affects the vector distance of the sample pairs for structure contrastive learning, a larger $\tau$ reduces the effect of structure contrastive learning, resulting in the reduction of ACC, NMI, and MF1 metrics for both datasets.

### v2

@@ III-F Model Training and Complexity Analysis
Input: i) a data graph $G$ with adjacency matrix $A$ and attribute matrix $X$, and ii) the number of training epochs $\varepsilon$
@@
// high-level graph preprocessing
2
[…]
4construct a dictionary $Dic$ by counting the number of specified patterns for each edge of $G$
[…]
// training process
6
7for epoch $\leftarrow 1,2,\cdots,\varepsilon$ do
8
9      $S^{\prime}\leftarrow W^{(2)}_{dnn}(\sigma(W^{(1)}_{dnn}S+b^{(1)}_{dnn}))+b^{(2)}_{dnn}$triangleright Eq. 2
10      $X^{\prime}\leftarrow W^{(2)}_{dnn}(\sigma(W^{(1)}_{dnn}X+b^{(1)}_{dnn}))+b^{(2)}_{dnn}$triangleright Eq. 3
11      $Z\leftarrow\mathbb{F}(A,AGG(S^{\prime},X^{\prime}))$triangleright Eq. 4
12      $Z^{H}\leftarrow\mathbb{F}(A^{H},AGG(S^{\prime},X^{\prime}))$triangleright Eq. 7
13      $\ell(z_{i}^{1}),\ell(z_{i}^{2})$ are obtained by calculating the similarity between positive samples and the similarity between negative samples with ($Z,Z^{H}$)
14      $\mathcal{L}\leftarrow\frac{1}{2N}\sum_{i=1}^{N}[\ell(z_{i}^{1})+\ell(z_{i}^{2})]$triangleright Eq. 11
15      update the GCLS2 network parameters by gradient descent to minimize $\mathcal{L}$
16
17return $Z$
@@
III-G Theoretical Analyses
In this section, we provide a theoretical justification for the GCLS2 model from Mutual Information (MI) perspective.
We find that the absolute value of our proposed loss Eq. 11 is a lower bound on the mutual information maximization between the node features and the embeddings of view 1 and view 2, which has been widely used in representation learning tasks [11, 28, 29, 30].
TABLE I: Statistics of the tested real-world graph datasets.
Datasets Type #Nodes #Edges Avg. Deg. #Attributes #Communities #Partitions
Cora Citation Network 2,708 5,429 4.01 1,433 7 1
Citeseer Citation Network 3,327 4,732 2.84 3,703 6 1
PubMed Citation Network 19,717 44,338 4.50 500 3 1
CoauthorCS Citation Network 18,333 81,894 8.93 6,805 15 1
AmazonPhotos Social Network 7,650 119,081 31.13 745 8 1
Email-Eu Social Network 1,005 25,571 50.93 NA 42 1
TWeibo Social Network 1,944,589 50,133,382 51.56 1,657 8 256
AmazonProducts Social Network 1,569,960 132,954,714 169.37 200 107 256
Lemma 1.
Let $\mathbf{X}$ be a variable with random uniform distribution after GNN coding, where $\mathbf{X}_{i}$ is the output vector of the node within the domain indicated by the GNN architecture. Given two embedding representations $\mathbf{Z}^{1},\mathbf{Z}^{2}\in\mathbb{R}^{L}$ of view 1 and view 2 are two random variables, their joint distribution is denoted as $p(\mathbf{Z}^{1},\mathbf{Z}^{2})$. Our objective $\mathcal{L}$ is a lower bound on the mutual information of $\mathbf{X}$ and $(\mathbf{Z}^{1},\mathbf{Z}^{2})$ as follows,
$-\mathcal{L}\leq I(\mathbf{X};\mathbf{Z}^{1},\mathbf{Z}^{2}).$ (15)
Proof:.
For a rigorous comparison, we compare our objective with the GRACE objective [31, 30, 11], which is defined under existing conditions as follows,
$\Gamma(z^{1}_{i})=\sum_{j=1}^{N}e^{sim(z_{i}^{1},z_{j}^{2})/\tau},$ (16)
$\gamma(z^{1}_{i})=\sum_{j=1,j\neq i}^{N}e^{sim(z_{i}^{1},z_{j}^{1})/\tau},$ (17)
$\small I_{ACE}(\mathbf{Z}^{1};\mathbf{Z}^{2})\triangleq\mathbb{E}_{\prod_{i}p(z_{i}^{1},z_{i}^{2})}[\frac{1}{N}\sum_{i=1}^{N}log\frac{e^{sim(z_{i}^{1},z_{i}^{2})/\tau}}{\Gamma(z^{1}_{i})+\gamma(z^{1}_{i})}],$ (18)
where $\prod_{i}p(z_{i}^{1},z_{i}^{2})$ denotes the joint distribution. To facilitate with the notational representation, we define $\Gamma(z^{1}_{i})=\sum_{j=1}^{N}e^{sim(z_{i}^{1},z_{j}^{2})/\tau}$, $\gamma(z^{1}_{i})=\sum_{j=1,j\neq i}^{N}e^{sim(z_{i}^{1},z_{j}^{1})/\tau}$. Please note that $\Gamma(z^{2}_{i})$ and $\gamma(z^{2}_{i})$ are defined symmetrically with $\Gamma(z^{1}_{i})$ and $\gamma(z^{1}_{i})$ respectively. So that our objective $\mathcal{L}$ can be rewritten as
$\small-\mathcal{L}=\mathbb{E}_{\prod_{i}p(z_{i}^{1},z_{i}^{2})}[\frac{1}{N}\sum_{i=1}^{N}log\frac{pos(z_{i}^{1})+e^{sim(z_{i}^{1},z_{i}^{2})/\tau}}{\sqrt{(\Gamma(z^{1}_{i})+\gamma(z^{1}_{i}))(\Gamma(z^{2}_{i})+\gamma(z^{2}_{i}))}}],$ (19)
where the $\sqrt{(\Gamma(z^{1}_{i})+\gamma(z^{1}_{i}))(\Gamma(z^{2}_{i})+\gamma(z^{2}_{i}))}$ is to illustrate the symmetry of $\Gamma(z^{1}_{i})+\gamma(z^{1}_{i})$ and $\Gamma(z^{2}_{i})+\gamma(z^{2}_{i})$, which are equal in value. Therefore, we can use the Eq. 18 Eq. 19 to deduce that
$\small\begin{split}-2\mathcal{L}&=I_{ACE}(\mathbf{Z}^{1};\mathbf{Z}^{2})-\mathbb{E}_{\prod_{i}p(z_{i}^{1},z_{i}^{2})}[\frac{1}{N}\sum_{i=1}^{N}log(1+\frac{pos(z_{i}^{1})}{e^{sim(z_{i}^{1},z_{i}^{2})/\tau}})]\\ &+I_{ACE}(\mathbf{Z}^{2};\mathbf{Z}^{1})-\mathbb{E}_{\prod_{i}p(z_{i}^{1},z_{i}^{2})}[\frac{1}{N}\sum_{i=1}^{N}log(1+\frac{pos(z_{i}^{2})}{e^{sim(z_{i}^{2},z_{i}^{1})/\tau}})]\\ &\leq I_{ACE}(\mathbf{Z}^{1};\mathbf{Z}^{2})+I_{ACE}(\mathbf{Z}^{2};\mathbf{Z}^{1}).\end{split}$ (20)
And according to [11, 30], the GRACE objective is a lower bound of the true MI, i.e.,
$I_{ACE}(\mathbf{Z}^{1};\mathbf{Z}^{2})\leq I(\mathbf{Z}^{1};\mathbf{Z}^{2}).$ (21)
Thus, we have the following inference
$-2\mathcal{L}\leq I(\mathbf{Z}^{1};\mathbf{Z}^{2})+\leq I(\mathbf{Z}^{2};\mathbf{Z}^{1})\leq 2I(\mathbf{Z}^{1};\mathbf{Z}^{2}).$ (22)
This is equivalent to
$-\mathcal{L}\leq I(\mathbf{Z}^{1};\mathbf{Z}^{2}).$ (23)
For three random variables $\mathbf{X}$, $\mathbf{Y}$, $\mathbf{Z}$ satisfying one Markov chain $\mathbf{X}\rightarrow\mathbf{Y}\rightarrow\mathbf{Z}$, we have $I(\mathbf{X};\mathbf{Z})\leq I(\mathbf{X};\mathbf{Y})$ holds. We find that $\mathbf{X}$, $\mathbf{Z}^{1}$, $\mathbf{Z}^{2}$ satisfy the chain relation $\mathbf{Z}^{1}\leftarrow\mathbf{X}\rightarrow\mathbf{Z}^{2}$. Actuality, $\mathbf{Z}^{1}$ and $\mathbf{Z}^{2}$ are conditionally independent after considering $\mathbf{X}$. This represents the existence of Markov chain $\mathbf{Z}^{1}\rightarrow\mathbf{X}\rightarrow\mathbf{Z}^{2}$, and we have $I(\mathbf{Z}^{1};\mathbf{Z}^{2})\leq I(\mathbf{Z}^{1};\mathbf{X})$. Since the chain relation $\mathbf{X}\rightarrow(\mathbf{Z}^{1},\mathbf{Z}^{2})\rightarrow\mathbf{Z}^{1}$ holds, it follows that $I(\mathbf{X};\mathbf{Z}^{1})\leq I(\mathbf{X};\mathbf{Z}^{1},\mathbf{Z}^{2})$. Combining the two inequalities above, we have the inequality
$I(\mathbf{Z}^{1};\mathbf{Z}^{2})\leq I(\mathbf{X};\mathbf{Z}^{1},\mathbf{Z}^{2}).$ (24)
According to Eq. 23 and Eq. 24, we can derive the final inequality as follows and complete the proof,
$-\mathcal{L}\leq I(\mathbf{X};\mathbf{Z}^{1},\mathbf{Z}^{2}).$ (25)
Analysis: Through Lemma 15, we can know that minimizing $\mathcal{L}$ is equivalent to maximizing the lower bound on the mutual information $I(\mathbf{X};\mathbf{Z}^{1},\mathbf{Z}^{2})$ between the input node features and the learned node representations, which theoretically provides the success of our model training. This is what the previous work [32, 33, 34, 35, 9] on contrastive learning has not focused on.
@@ III-G Theoretical Analyses
IV Experimental Evaluation
In this section, we report the effectiveness and efficiency of our proposed GCLS2 approach on real-world graph datasets by answering the following four research questions (RQs).
@@ IV Experimental Evaluation
  • RQ2 (Applicability of GCLS2): What are impacts of different parameter settings (e.g., semantic dimension $d$ and temperature parameter $\tau$) on the GCLS2 performance?
  • RQ3 (Scalability of GCLS2): Can GCLS2+HGP algorithm perform the community detection efficiently on large-scale graphs?
  • RQ4 (Benefits of GCLS2): Can GCLS2 significantly improve the accuracy, efficiency, and modularity of the community detection?
IV-A Experimental Settings
Datasets and Baselines: We use six public real-world graph datasets: the widely-used citation networks (i.e., Cora, Citeseer, and PubMed [36]), the co-authorship network and the product co-purchase network (i.e., CoauthorCS and AmazonPhotos [37], resp.), and the e-mail network between members of a European research institution (Email_Eu [38]) to evaluate the effectiveness of our GCLS2 method on the community detection task. We also use two large-scale public real-world graphs, TWeibo [39] (i.e., a social network of Tencent Weibo users) and AmazonProducts [40] (i.e., a product review network of Amazon), to demonstrate the applicability and effectiveness of our proposed high-level graph partitioning algorithms HGP and GCLS2. The details of the datasets are in Table I. We use the notation “NA” to denote the absence of attributes for each node in the graph dataset. We compare our approach with two supervised learning baselines (i.e., GCN [41] and GAT [42]) and seven unsupervised learning baselines, including DGI [43], MVGRL [44], GRACE [11], GCA [12], SUGRL [45], NCLA [34], and ASP [35]. We compare the HGP in Algorithm 1 with three classical graph partitioning algorithms (i.e., METIS [22], LGD [23], and Hash [46] algorithms) on large-scale graphs, in terms of the efficiency of graph partitioning and the accuracy of the community detection.
Metrics: We use classical community detection metrics, normalized mutual information (NMI), accuracy (ACC), and Macro-F1 score (MF1) [47], to evaluate the effectiveness of the model detection. For all the above metrics, the higher their values, the better. To evaluate the efficiency of GCLS2 and HGP, we measured the wall clock times of online community detection and graph partitioning, respectively. We evaluate the load balancing capability of graph partitioning algorithms by calculating the maximum load difference between partitions; the smaller the maximum load difference, the better the load balancing capability.
TABLE IV: Performance of GCLS2 on large-scale graph datasets.
Methods TWeibo AmazonProducts
ACC NMI MF1 ACC NMI MF1
GCN 33.34 20.56 38.39 30.52 23.75 38.12
GAT 35.78 24.26 36.91 41.72 36.87 43.57
MVGRL 38.68 30.47 37.77 50.32 43.48 51.46
GRACE 35.20 28.92 26.19 46.05 40.94 49.31
SUGRL 39.57 34.25 45.79 54.93 47.82 55.68
GCLS2 42.94 44.72 59.92 61.79 53.01 63.54
Figure 5: The community detection inference time comparison of different methods on large-scale graph datasets.
@@ IV-A Experimental Settings
We run all the experiments on a PC with an NVIDIA RTX 4090 GPU and 24GB of memory. All algorithms are implemented in Python and executed with the Python 3.11 interpreter.
TABLE V: Performance of the variants on Cora and Email-Eu.
@@
(a) partitioning efficiency
(b) load difference
(c) analysis of p-parameters
Figure 6: Performance evaluation of the HGP algorithm.
IV-B The Performance Evaluation – RQ1
To answer RQ1, we evaluate the performance of the baselines and GCLS2 with and without attributes. For the without-attribute case, we test by replacing the attribute matrix with an adjacency matrix. Table II shows the performance of the baselines and GCLS2 on different graph networks with attributes. Table III shows the performance without attributes. Overall, the tables show that our GCLS2 performs strongly on all six datasets, especially in the case of without attributes, where it has quite an advantage. These performances validate the excellence of our GCLS2 method. Specifically, we have the following observations.
First, GCLS2 has an advantage in comparison to GCN, which we believe is due to the fact that GCLS2 deeply mines the structure information of the community and enhances the representation of this information through structure contrastive learning, and also compared to GCN, which has an average reduction of 4.76% in the case of without attributes, GCLS2 has a reduction of 3.05%, where excluding the Citeseer dataset, it has only a reduction of 1%. This emphasizes the importance of considering structure semantics in community detection.
Second, we find that the node-graph level contrastive learning methods (i.e., DGI, MVGRL) are significantly more effective than the node-node level (i.e., GRACE, NCLA, GCLS2) in the CoauthorCS dataset. We believe that the CoauthorCS with a large number of attributes (i.e., 6,805) is represented on the data graph with rich information, which is more sensitive to the community information, and at the same time, removing the attribute information, the node-graph level effect is not outstanding, which confirms our thoughts. In contrast, our proposed GCLS2, which also focuses on structure semantic representation at the node-node level, has the highest detection performance without attribute information.
Due to the memory limitations of the experimental machines, most network models, with their complex structures, require significant memory resources during online training. We successfully evaluate the performance of GCN, GAT, MVGRL, GRACE, and SUGRL with our proposed HGP partitioning algorithm with partition size $p=256$ on two large-scale datasets: TWeibo and AmazonProducts. As shown in Table IV, our GCLS2 method achieves the best performance in terms of accuracy and modularity compared to baseline methods. On the TWeibo dataset, GCLS2 outperforms the second-best method by 3% in accuracy and 10% in modularity. On the dense AmazonProducts dataset, it exceeds the second-best by 7% in accuracy, highlighting GCLS2’s advantage in handling dense graphs. Moreover, GCLS2 not only outperforms supervised learning methods such as GCN and GAT, but also learns similar embeddings with potential community structures by high-level structural contrastive learning. This approach overcomes the limitation of traditional methods, which focus on aligning embeddings from different views of the same vertex, and proves effective for community detection on dense and large-scale datasets.
To further validate the computational cost advantage of our GCLS2 model, we measure the inference time of the model on different large-scale datasets. Figure 5 shows the inference time of our model and baselines on two large-scale datasets (i.e., TWeibo and AmazonProducts). The results show that GCLS2 achieves the shortest inference time among all GCL models. This is primarily because GCLS2 avoids additional data augmentation or node feature resampling. For GAT, the complex attention mechanism increases the computational cost of the model. Most GCL models are designed based on GCNs, making a single GCN the fastest in terms of inference. In comparison, GRACE generates multiple views by removing edges and masking node features; MVGRL not only generates multiple views but also uses a discriminator to measure agreement across different views; and SUGRL further samples node features. These operations increase inference costs. In contrast, GCLS2 uses lightweight SSS feature extraction and structure loss calculation, with no additional overhead, resulting in faster inference.
IV-C The Ablation Study – RQ2
To answer RQ2, we study the variants of the proposed GCLS2 method and analyze the performance of each of the important modules of Figure 2. Specifically, we have three variants: i) we separate the structure similarity matrix $S$, which represents dense community information, as a variant, i.e., GCLS2 w/o $S$; ii) we separate the structure similarity semantic (SSS) module, which extracts low-level semantic representations as a variant, i.e., GCLS2 w/o SSS; iii) we separate the structure contrastive loss that reinforces the expression of community structure information as a variant, i.e., GCLS2 w/o SCL, and the rest of the processing for each variant was the same as for GCLS2. Table V shows the performance of the GCLS2’s variants on Cora and Email-Eu datasets. From Table V, the GCLS2 consistently has the highest accuracy among all the variants on the two datasets. Compared to the three variants, the GCLS2 achieves significant gains on Email-Eu and Cora. Especially in the Email-Eu, the accuracy of using SCL increased by 8%. This reflects that the structure semantic information is effective in community detection, and the structure contrastive learning has shown good performance improvement on Email-Eu.
@@ IV-C The Ablation Study – RQ2
IV-D The HGP Algorithm Evaluation – RQ3
We evaluate the performance of the GCLS2 combined with the HGP algorithm on large-scale datasets and compare the HGP algorithm with state-of-the-art graph partitioning algorithms, including METIS, LDG, and Hash. As shown in Table VI, the GCLS2 with HGP achieves the best performance on the TWeibo and AmazonProducts datasets, outperforming the second-best by 4% in MF1. Note that “OOM” denotes out-of-memory errors. Figure 6(b) shows that the maximum load difference of LDG on TWeibo and AmazonProducts is close to $10^{6}$, illustrating the inability of the LDG algorithm to maintain good load balancing. This makes it challenging for small-memory machines to support online training.
Figure 6 presents the performance evaluation of the HGP algorithm. It can be seen that the HGP algorithm can partition hundreds of millions of edge-level graphs in approximately $10^{3}$ seconds while ensuring load balancing across 256 partitions. In contrast, the Hash algorithm completes partitioning in less than 1 second by performing only a random hash computation for each node. Moreover, as seen in Figure 6(b), the HGP algorithm achieves good load balancing, reducing the maximum load difference by one order of magnitude compared to METIS and three orders of magnitude compared to LDG. From Figure 6(c), when the partition size $p=32$, the HGP algorithm completes the partitioning of graphs with tens of millions of edges in just 155.91 seconds. Combining these findings with Table IV, it is clear that the GCLS2+HGP method performs well in terms of efficiency and accuracy of community detection in large-scale datasets.
TABLE VI: Performance of GCLS2 with different graph partitioning algorithms on large-scale graph datasets.
Methods TWeibo AmazonProducts
ACC NMI MF1 ACC NMI MF1
GCLS2+METIS 41.83 42.90 55.31 58.73 51.82 59.40
GCLS2+LDG OOM OOM OOM OOM OOM OOM
GCLS2+Hash 38.62 35.33 40.15 55.52 49.29 57.35
GCLS2+HGP 42.94 44.72 59.92 61.79 53.01 63.54